# Preserving Data Secrecy in Decentralized Inter-organizational Process Mining

### 05 Jan 2024 v2

@@ Abstract
Collaborative information systems architectures Trusted execution environments Confidential computing1 Introduction
@@
In today’s business landscape, organizations constantly seek ways to enhance operational efficiency, increase performance, and gain valuable insights to improve their processes. Process mining offers techniques to discover, monitor, and improve business processes by extracting knowledge from chronological records known as event logs [4]. Process-aware information systems record events referring to activities and interactions within a business process. The vast majority of process mining contributions consider intra-organizational settings, in which business processes are executed inside individual organizations. However, organizations increasingly recognize the value of collaboration and synergy in achieving operational excellence. Inter-organizational business processes involve several independent organizations cooperating to achieve a shared objective [2]. Despite the advantages of transparency, performance optimization, and benchmarking that companies can gain, inter-organizational process mining raises challenges that hinder its application. The major issue concerns confidentiality. Companies are reluctant to share private information required to execute process mining algorithms with their partners [23]. Indeed, letting sensitive operational data traverse organizational boundaries introduces concerns about data privacy, security, and compliance with internal regulations [25]. Trusted Execution Environments (TEEs) [28] can serve as fundamental enablers to balance the need for insights with the need to protect sensitive information in inter-organizational settings. TEEs offer secure contexts that guarantee code integrity and data confidentiality before, during, and after its utilization.
@@
In this paper, we propose CONFINE, a novel approach and tool aimed at enhancing collaborative information system architectures with secrecy-preserving process mining capabilities in a decentralized fashion. It resorts to trusted applications running in TEEs to preserve the secrecy and integrity of shared data. To pursue this aim, we design a decentralized architecture for a four-staged protocol: (i) The initial exchange of preliminary metadata, (ii) the attestation of the miner entity, (iii) the secure transmission and privacy-preserving merge of encrypted information segments amid multiple parties, (iv) the isolated and verifiable computation of process discovery algorithms on joined data. We evaluate our proof-of-concept implementation against synthetic and real-world-based data with a convergence test followed by experiments to assess the scalability of our approach.
@@ 2 Related Work
Despite the relative recency of this research branch across process mining and collaborative information systems, scientific literature already includes noticeable contributions to inter-organizational process mining. The work of Müller et al. [26] focuses on data privacy and security within third-party systems that mine data generated from external providers on demand. To safeguard the integrity of data earmarked for mining purposes, their research introduces a conceptual architecture that entails the execution of process mining algorithms within a cloud service environment, fortified with Trusted Execution Environments. Drawing inspiration from this foundational contribution, our research work seeks to design a decentralized approach characterized by organizational autonomy in the execution of process mining algorithms, devoid of synchronization mechanisms involvement taking place between the involved parties. A notable departure from the framework of Müller et al. lies in the fact that here each participating organization retains the discretion to choose when and how mining operations are conducted. Moreover, we bypass the idea of fixed roles, engineering a peer-to-peer scenario in which organizations can simultaneously be data provisioners or miners. Elkoumy et al. [17, 16] present Shareprom. Like our work, their solution offers a means for independent entities to execute process mining algorithms in inter-organizational settings while safeguarding their proprietary input data from exposure to external parties operating within the same context. Shareprom’s functionality, though, is confined to the execution of operations involving event log abstractions [3] represented as directed acyclic graphs, which the parties employ as intermediate pre-elaboration to be fed into secure multiparty computation (SMPC) [13]. As the authors remark, relying on this specific graph representation imposes constraints that may prove limiting in various process mining scenarios. In contrast, our approach allows for the secure, ciphered transmission of event logs to process mining nodes as a whole. Moreover, SMPC-based solutions require computationally intensive operations and synchronous cooperation among multiple parties, which make these protocols challenging to manage as the number of participants scales up [32]. In our research work, individual computing nodes run the calculations, thus not requiring synchronization with other machines once the input data is loaded.
@@ 3 Motivating Scenario
For our motivating scenario, we focus on a simplified hospitalization process for the treatment of rare diseases. The process model is depicted as a BPMN diagram in Fig. 1 and involves the cooperation of three parties: a hospital, a pharmaceutical company, and a specialized clinic. For the sake of simplicity, we describe the process through two cases, recorded by the information systems as in Table 2. Alice’s journey (case 312) begins when she enters the hospital for the preliminary examinations (patient hospitalized, PH). The hospital then places an order for the drugs (OD) to the pharmaceutical company for treating Alice’s specific condition. Afterwards, the pharmaceutical company acknowledges that the drugs order is received (DOR), proceeds to produce the drugs in the laboratory (PDL), and ships the drugs (SD) back to the hospital. Upon receiving the medications, the hospital administers the drug (AD), and conducts an assessment to determine if Alice can be treated internally. If specialized care is required, the hospital transfers the patient (TP) to the specialized clinic. When the patient arrives from the hospital (PAFH), the specialized clinic performs in-depth analyses (PIA) and proceeds with the treatment (PT). Once the specialized clinic had completed the evaluations and verified the response to the treatment (VRT), it transfers the patient back (TPB). The hospital receives the patient back (RPB) and prepares the clinical documentation (PCD). If Alice has successfully recovered, the hospital declares the patient as healed (DPH). When Alice’s treatment is complete, the hospital discharges the patient (DP). Bob (case 711) enters the hospital a few hours later than Alice. His hospitalization process is similar to Alice’s. However, he does not need specialized care, and his case is only treated by the hospital. Therefore, the hospital performs the response to treatment analyses (PRTA) instead of transferring him to the specialized clinic.
@@
Both the National Institute of Statistics of the country in which the three organizations reside and the University that hosts the hospital wish to uncover information on this inter-organizational process for reporting and auditing purposes [19] via process analytics. The involved organizations share the urge for such an analysis and wish to be able to repeat the mining task also in-house. The hospital, the specialized clinic, and the pharmaceutical company have a partial view of the overall unfolding of the inter-organizational process as they record the events stemming from the parts of their pertinence. In Table 1, we show the cases 312 and 711 and the corresponding traces recorded by the hospital (i.e., and ), the specialized clinic (i.e., and ), and the pharmaceutical company (i.e., and ). Those traces are projections of the two combined ones for the whole inter-organizational process: ⟨PH, COPA, OD, DOR, PDL, SD, RD, AD, TP, PAFH, PIA, PT, VRT, TPB, RPB, DPH, PCD, DP⟩ and ⟨PH, COPA, OD, DOR, PDL, SD, RD, AD, TP, DPH, PCD, DP⟩. Results stemming from the analysis of the local cases would not provide a full picture. Data should be merged. However, to preserve the privacy of the people involved and safeguard the confidentiality of the information, the involved parties cannot give open access to their traces to other organizations. The diverging interests (being able to conduct process mining on data from multiple sources without giving away the local event logs in-clear) motivate our research. In the following, we describe the design of our solution.
@@ 4 Design
In this section, we present the high-level architecture of the CONFINE framework. We consider the main functionalities of each component, avoiding details on the employed technologies discussed in the next sections.
@@
The CONFINE architecture at large. Our architecture involves different information systems running on multiple machines. An organization can take at least one of the following roles: provisioning if it delivers local event logs to be collaboratively mined; mining if it applies process mining algorithms using event logs retrieved from provisioners. In Fig. 2, we propose the high-level schematization of the CONFINE framework. In our solution, every organization hosts one or more nodes hosting components (the names of which will henceforth be formatted with a teletype font). Depending on the played role, nodes come endowed with a Provisioner or a Secure Miner component, or both. The Provisioner component consists of the following two main sub-components. The Log Recorder registers the events taking place in the organizations’ systems. The Log Provider delivers on-demand data to mining players. The hospital and all other parties in our example record Alice and Bob’s cases using the Log Recorder. The Log Recorder is queried by the Log Provider for event logs to be made available for mining. The latter controls access to local event logs by authenticating data requests by miners and rejecting those that come from unauthorized parties. In our motivating scenario, the specialized clinic, the pharmaceutical company, and the hospital leverage Log Providers to authenticate the miner party before sending their logs. The Secure Miner component shelters external event logs inside a protected environment to preserve data confidentiality and integrity. Notice that Log Providers accept requests issued solely by Secure Miners. Next, we provide an in-depth focus on the latter.
The Secure Miner. The primary objective of the Secure Miner is to allow miners to securely execute process mining algorithms using event logs retrieved from provisioners such as the specialized clinic, pharmaceutical company, and the hospital in our example. Secure Miners are isolated components that guarantee data inalterability and confidentiality. In Fig. 3, we show a schematization of the Secure Miner, which consists of four sub-components: (i) Log Requester; (ii) Log Receiver; (iii) Log Manager; (iv) Log Elaborator. The Log Requester and the Log Receiver are the sub-components that we employ during the event log retrieval. Log Requesters send authenticable data requests to the Log Providers. The Log Receiver collects event logs sent by Log Providers and entrusts them to the Log Manager, securing them from accesses that are external to the Secure Miner. Miners of our motivating scenario, such as the university and the national institute of statistics, employ these three components to retrieve and store Alice and Bob’s data. The Log Elaborator merges the event data locked in the Secure Miner to have a global view of the inter-organizational process comprehensive of activities executed by each involved party. Thereupon, it executes process mining algorithms in a protected environment, inaccessible from the outside computation environment. In our motivating scenario, the Log Elaborator combines the traces associated with the cases of Alice (i.e., , , and ) and Bob (i.e, , , and ), generates the chronologically sorted traces and , and feeds them into the mining algorithms (see the bottom-right quadrant of Table 1).
@@ 5 Realization
In this section, we outline the technical aspects concerning the realization of our solution. Therefore, we first present the enabler technologies through which we instantiate the design principles presented in Sect. 4. After that, we discuss the CONFINE interaction protocol. Finally, we show the implementation details.
@@ 5.1 Deployment
Provisioner Nodes host the Provisioner’s components, encompassing the Log Recorder and the Log Provider. The Process-Aware Information System (PAIS) manifests the Log Recorder [15]. The PAIS grants access to the Log Server, enabling it to retrieve event log data. The Log Server, on the other hand, embodies the functionalities of the Log Provider, implementing web services aimed at handling remote data requests and providing event log data to miners.
The Miner Node is characterized by two distinct execution environments: the Operating System (OS) and the Trusted Execution Environment (TEE) [28]. TEEs establish isolated contexts separate from the OS, safeguarding code and data through hardware-based encryption mechanisms. This technology relies on dedicated sections of a CPU capable of handling encrypted data within a reserved section of the main memory [12]. By enforcing memory access restrictions, TEEs aim to prevent one application from reading or altering the memory space of another, thus enhancing the overall security of the system. This dedicated areas in memory are, however, limited. Once the limits are exceeded, TEEs have to scout around in outer memory areas, thus conceding the opportunity to malicious reader to understand the saved data based on the memory reads and writes. To avoid this risk, TEE implementations often raise errors that halt the program execution when the memory demand goes beyond the available space. Therefore, the design of secure systems that resort to TEEs must take into account that memory consumption must be kept under control. We leverage the security guarantees provided by TEEs [20] to protect a Trusted App responsible for fulfilling the functions of the Secure Miner and its associated sub-components. The TEE ensures the integrity of the Trusted App code, protecting it against potential malicious manipulations and unauthorized access by programs running within the Operating System. Additionally, we utilize the isolated environment of TEEs to securely store event log data (e.g., Alice and Bob’s cases). The TEE retains a private key in the externally inaccessible section of memory, paired with a public key in a Rivest-Shamir-Adleman (RSA) [27] scheme for attestation (only the owner of the private key can sign messages that are verifiable via the public key) and secure message encryption (only the owner of the private key can decode messages that are encrypted with the corresponding public key). In our solution, access to data located in the TEE is restricted solely to the Trusted App. Users interact with the Trusted App through the Trusted App Interface, which serves as the exclusive communication channel. The Trusted App offers secure methods, invoked by the Trusted App Interface, for safely receiving information from the Operating System and outsourcing the results of computations.

### 09 Apr 2024 v3

# Trusted Execution Environment for Decentralized Process Mining

@@ Abstract
Collaborative information systems architectures Inter-organizational process mining TEE Confidential computing1 Introduction
@@
In today’s business landscape, organizations constantly seek ways to enhance operational efficiency, increase performance, and gain valuable insights to improve their processes. Process mining offers techniques to discover, monitor, and improve business processes by extracting knowledge from chronological records recorded by process-aware information systems, i.e., the event logs [13]. The vast majority of process mining contributions consider intra-organizational settings, in which processes are executed inside individual organizations. However, organizations increasingly recognize the value of collaboration and synergy in achieving operational excellence. Inter-organizational business processes involve several independent organizations cooperating to achieve a shared objective. Process mining can bring the advantages of transparency, performance optimization, and benchmarking in this context [2]. Since different process data owners feed separate mining nodes, this setting characterizes what we call decentralized process mining.
Companies, though, are reluctant to share private information required to execute process mining algorithms with external parties [25], thus hindering its adoption. Letting sensitive operational data traverse organizational boundaries introduces concerns about data secrecy, security, and compliance with internal regulations [27]. To address this issue, the majority of research endeavors have focused thus far on the alteration of input data or of intermediate analysis by-products, with the aim to impede the counterparts from reconstructing the original information sources [18, 17, 16, 15]. These preemptive solutions have the remarkable merit of neutralizing information leakage by malicious parties a priori. Nevertheless, they entail an ex-ante information loss, thus compromising downstream process mining capabilities [18, 17], or require the execution of computationally heavy protocols undermining scalability [16, 15].
To overcome these limitations, we propose CONFINE, a novel approach and tool aimed at enhancing collaborative information system architectures with secrecy-preserving process mining capabilities. To secure information secrecy during the exchange and elaboration of data, our solution resorts to Trusted Execution Environments (TEEs) [30], namely hardware-secured contexts that guarantee code integrity and data confidentiality before, during, and after their utilization. Owing to these characteristics, CONFINE lets information be securely transferred beyond the organizations’ borders. Therefore, computing nodes other than the information provisioners can aggregate and elaborate the original, unaltered process data in a secure, externally inaccessible vault. Also, CONFINE is capable of providing these guarantees while demanding scalable computational overhead.
@@
The decentralized architecture of CONFINE supports a four-staged protocol: (i) The initial exchange of preliminary metadata, (ii) the attestation of the mining entities, (iii) the secure transmission and secrecy-preserving merge of encrypted information segments amid multiple parties, (iv) the isolated and verifiable computation of process discovery algorithms on joined data. We evaluate our proof-of-concept implementation against synthetic and real-world data with a convergence test followed by experiments to assess the scalability of our approach. Since TEEs operate with dedicated memory pages shielded from access by external entities (operating system included), thus entailing a hardware constraint on computation space, we endow our experiments with an analysis of memory usage, too.
@@ 2 Related Work
The scientific literature already includes noticeable contributions to process mining in a decentralized setting with a focus on data secrecy, despite the relative recency of this research branch across process mining and collaborative information systems. The work of Müller et al. [28] revolves around data privacy and security within third-party systems that mine data generated from external providers on demand. To safeguard the integrity of data earmarked for mining purposes, their research introduces a conceptual architecture that entails the execution of process mining algorithms within a cloud service environment, fortified with Trusted Execution Environments. Drawing inspiration from this foundational contribution, our research work seeks to design a decentralized approach characterized by organizational autonomy in the execution of process mining algorithms, devoid of synchronization mechanisms taking place between the involved parties. A notable departure from the framework of Müller et al. lies in the fact that here each participating organization retains the discretion to choose when and how mining operations are conducted. Moreover, we bypass the idea of fixed roles, engineering a peer-to-peer scenario in which organizations can simultaneously be data provisioners or miners. Fahrenkrog-Petersen et al. [18, 17] theorize the PRETSA algorithms family, namely a set of event log sanitization techniques that perform step-wise transformations of prefix-tree event log representation into a sanitized output ensuring k-anonimization and t-closeness. While these algorithms effectively minimize information loss, they introduce targeted approximations within the original event log, which may compromise the exactness of process mining results or inhibit mining tasks. In contrast, our research proposes an architecture wherein secure computational vaults collect event logs devoid of upstream alterations and protect them at runtime, thus generating results derived directly from the original information source. Elkoumy et al. [16, 15] present Shareprom. Like our work, their solution offers a means for independent entities to execute process mining algorithms in inter-organizational settings while safeguarding the proprietary input data from exposure to external parties operating within the same context. Shareprom’s functionality, though, is confined to the execution of operations involving event log abstractions [3] represented as directed acyclic graphs, which the parties employ as intermediate pre-elaboration to be fed into secure multiparty computation (SMPC) [12]. As the authors remark, relying on this specific graph representation imposes constraints that may turn out to be limiting in a number of process mining scenarios. In contrast, our approach allows for the secure, ciphered transmission of event logs (or segments thereof) to process mining nodes. Moreover, SMPC-based solutions require computationally intensive operations and synchronous cooperation among multiple parties, which make these protocols challenging to manage as the number of participants scales up [37]. In our research work, individual computing nodes run the calculations, thus not requiring synchronization with other machines once the input data is loaded.
@@ 3 Motivating Scenario
To provide a running example and motivating scenario for our investigation, we focus on a simplified hospitalization process for the treatment of rare diseases. The process model is depicted as a BPMN diagram in Fig. 1 and involves the cooperation of three parties: a hospital, a pharmaceutical company, and a specialized clinic. For the sake of simplicity, we describe the process through two cases, recorded by the information systems as in Table 1. Alice’s journey (case 312) begins when she enters the hospital for the preliminary examinations (patient hospitalized, PH). The hospital then places an order for the drugs (OD) to the pharmaceutical company for treating Alice’s specific condition. Afterwards, the pharmaceutical company acknowledges that the drugs order is received (DOR), proceeds to produce the drugs in the laboratory (PDL), and ships the drugs (SD) back to the hospital. Upon receiving the medications, the hospital administers the drug (AD), and conducts an assessment to determine if Alice can be treated internally. If specialized care is required, the hospital transfers the patient (TP) to the specialized clinic. When the patient arrives from the hospital (PAFH), the specialized clinic performs in-depth analyses (PIA) and proceeds with the treatment (PT). Once the specialized clinic had completed the evaluations and verified the response to the treatment (VRT), it transfers the patient back (TPB). The hospital receives the patient back (RPB) and prepares the clinical documentation (PCD). If Alice has successfully recovered, the hospital declares the patient as healed (DPH). When Alice’s treatment is complete, the hospital discharges the patient (DP). Bob (case 711) enters the hospital a few hours later. His hospitalization process is similar to Alice’s. However, he does not need specialized care, and his case is only treated by the hospital. Therefore, the hospital performs the response to treatment analyses (PRTA) instead of transferring him to the specialized clinic.
@@
Both the National Institute of Statistics of the country in which the three organizations reside and the University that hosts the hospital wish to uncover information on this inter-organizational process for reporting and auditing purposes via process analytics [21]. The involved organizations share the urge for such an analysis and wish to be able to repeat the mining task also in-house. The hospital, the specialized clinic, and the pharmaceutical company have a partial view of the overall unfolding of the inter-organizational process as they record the events stemming from the parts of their pertinence. In Table 1, we show cases 312 and 711 and the corresponding traces recorded by the hospital (i.e., and ), the specialized clinic (i.e., and ), and the pharmaceutical company (i.e., and ). Those traces are projections of the two combined ones for the whole inter-organizational process: ⟨PH, COPA, OD, DOR, PDL, SD, RD, AD, TP, PAFH, PIA, PT, VRT, TPB, RPB, DPH, PCD, DP⟩ and ⟨PH, COPA, OD, DOR, PDL, SD, RD, AD, PRTA, PCD, DPH, DP⟩. Results stemming from the analysis of the local cases would not provide a full picture. Data should be merged. However, to safeguard the confidentiality of the information, the involved parties cannot give other organizations open access to their traces. The diverging interests (being able to conduct process mining on data from multiple sources without giving away the local event logs in-clear) motivate our research.
We remark that the problem we aim to solve spans across an array of domains beyond healthcare. It particularly applies to scenarios in which one or more parties are interested in process analytics outcomes based on data they bear but cannot be disclosed to the other process actors or to the miners. In the supply chain realm, e.g., the extraction of aggregate knowledge about trends and management guidelines is called for, but the acquisition of competitive advantage out of knowledge leakage must be prevented [33]. In personal informatics, company-wide work routine monitoring and analysis are desirable, though the details of individual participants should be sheltered from inquisitive inspections [31].
@@ 4 Design
Our goal is to enable the secure aggregation and elaboration of original, unaltered event logs from decentralized sources in dedicated environments that potentially lie beyond the individual organizations’ information perimeter. With this objective in mind, we devise the Secure Miner component, which is capable of safeguarding data merge and processing by running certified code in an isolated execution vault. Thus, we decouple provisioning from treatment, and the two tasks can be carried out by distinct computing nodes. Here, we introduce CONFINE’s key components, with a special focus on the Secure Miner.
The CONFINE architecture at large. Our architecture involves different information systems running on multiple machines. An organization can take at least one of the following roles: provisioning if it delivers local event logs to be collaboratively mined; mining if it applies process mining algorithms using event logs retrieved from provisioners. Figure 2 depicts the high-level schematization of the CONFINE framework. In our solution, each organization hosts one or more nodes encompassing diverse components (the names of which will henceforth be formatted with a teletype font). Depending on the played role, nodes come endowed with a Provisioner or a Secure Miner, or both. The Provisioner component, in turn, consists of the following two sub-components. The Log Recorder registers the events taking place in the organizations’ systems. The Log Provider delivers on-demand data to miners. The hospital and all other parties in our example record Alice and Bob’s cases using the Log Recorder. The Log Recorder, in turn, is queried by the Log Provider for event logs to be made available for mining. The latter controls access to local event logs by authenticating data requests by miners and rejecting those that come from unauthorized parties. In our motivating scenario, the specialized clinic, the pharmaceutical company, and the hospital leverage Log Providers to authenticate the miner before sending their logs. The Secure Miner component shelters external event logs inside a protected environment to preserve data confidentiality and integrity. Notice that Log Providers accept requests issued solely by Secure Miners. Next, we provide an in-depth focus on the latter.
The Secure Miner. The primary objective of the Secure Miner is to allow miners to securely execute process mining algorithms using event logs retrieved from provisioners (the specialized clinic, the pharmaceutical company, and the hospital in our example). Secure Miners are isolated components that guarantee data inalterability and confidentiality.
@@
Figure 3 illustrates a schematization of the Secure Miner, which consists of four sub-components: (i) the Log Requester; (ii) the Log Receiver; (iii) the Log Manager; (iv) the Log Elaborator. The Log Requester and the Log Receiver are the sub-components that we employ during the event log retrieval. Log Requesters send authenticable data requests to the Log Providers. The Log Receiver collects event logs sent by Log Providers and entrusts them to the Log Manager, securing them from accesses that are external to the Secure Miner. Miners of our motivating scenario, such as the university and the national institute of statistics, employ these three components to retrieve and store Alice and Bob’s data. The Log Manager merges the event data locked in the Secure Miner to have a global view of the inter-organizational process comprehensive of activities executed by each involved party. The Log Elaborator executes process mining algorithms in a protected environment, inaccessible from the outside computation environment. In our motivating scenario, the Log Manager combines the traces associated with the cases of Alice (i.e., , , and ) and Bob (i.e, , , and ), generates the chronologically sorted traces and , and feeds them into the Log Elaborator’s mining algorithms (see the bottom-right quadrant of Table 1).
@@ 5 Realization
Thus far, we have outlined the main functionalities of each component at large. Here we discuss the technical aspects concerning the realization of our solution. We first present the technologies through which we enable the design principles in Sect. 4. Then, we discuss the CONFINE interaction protocol. Finally, we show the implementation details.
@@ 5.1 Deployment
Provisioner Nodes host the Provisioner’s components, i.e., the Log Recorder and the Log Provider. The Process-Aware Information System (PAIS) manifests the Log Recorder [14]. The PAIS grants access to the Log Server, enabling it to retrieve event log data. The Log Server, on the other hand, embodies the functionalities of the Log Provider, implementing services that handle remote data requests and provide event log data to the miners. The Miner Node is characterized by two distinct execution environments: the Operating System (OS) and the Trusted Execution Environment (TEE) [30]. TEEs establish isolated contexts separate from the OS, safeguarding code and data through hardware-based encryption mechanisms. This technology relies on dedicated sections of a CPU that handle encrypted data within a reserved section of the main memory [11].
By enforcing memory access restrictions, TEEs aim to prevent one application from reading or altering the memory space of another, thus enhancing system security. These dedicated areas in memory are limited, though. Once the limits are exceeded, TEEs have to scout around in outer memory areas, thus conceding the opportunity to malicious readers to understand the saved data based on the reads and writes. To avoid this risk, TEE implementations often raise errors that halt the program execution when memory demand goes beyond the available space. Therefore, the design of secure systems that resort to TEEs must take into account that memory consumption must be kept under control.
We leverage the security guarantees provided by TEEs [22] to protect a Trusted App responsible for fulfilling the functions of the Secure Miner and its associated sub-components. Our TEE component ensures the integrity of the Trusted App code, protecting it against potential malicious manipulations and unauthorized access by programs running within the OS. Additionally, we utilize the isolated environment of the TEE to securely store event log data (e.g., Alice and Bob’s cases). The TEE retains a private key in its inaccessible memory section, paired with a public key in a Rivest-Shamir-Adleman (RSA) [29] scheme for attestation (only the owner of the private key can sign messages in a way that is verifiable via the public key) and secure message encryption (only the owner of the private key can decode messages that are encrypted with the corresponding public key). In our solution, access to data located in the TEE is restricted to the sole Trusted App. Users interact with the Trusted App through the TEE Interface, which serves as the exclusive communication channel. The Trusted App offers secure methods, invoked by the Trusted App Interface, for safely receiving information from the OS and outsourcing the results of computations.